# Joint Precoding and Phase-Shift Optimization for Beyond-Diagonal RIS-Aided ISAC System

### 10 Mar 2026 v1

@@ Abstract
Beyond diagonal reconfigurable intelligent surfaces (BD-RIS) can realize the interconnection between reflecting elements through the impedance network, thereby providing a new approach for the performance improvement of integrated sensing and communication (ISAC) systems. This paper investigates the optimization problem of BD-RIS-aided multiuser ISAC system, aiming to achieve the flexible design of trade-offs between communication and sensing performance. Specifically, we propose an optimization framework jointly combining the multiuser interference management and sensing beam gain approximation method. By jointly optimizing the precoding vector and RIS phase-shift matrix, improving the multiuser communication sum rate through the proposed interference management method, and enhancing the system sensing performance through the beam gain approximation method. For the resulting non-convex weighted optimization problem, we employ the alternating optimization (AO) algorithm to decouple it into two subproblems of precoding vector and phase-shift matrix optimization, with each step admitting closed-form solutions. Simulation results demonstrate that the proposed BD-RIS-aided ISAC system can achieve significant improvement in the trade-offs between communication and sensing performance than the traditional diagonal RIS, verifying the effectiveness of the proposed optimization framework.
@@ I Introduction
The wireless propagation environment has a significant impact on the performance of ISAC systems. Severe occlusion in scenarios such as dense urban areas often leads to the absence of line-of-sight paths, which causes degradation in communication quality and limited sensing accuracy [5]. Reconfigurable intelligent surface (RIS), as an emerging low-cost technology, provides a feasible solution to mitigate this problem. By dynamically adjusting the phase-shifts of reflecting elements, RIS can reconfigure wireless channels, thereby enhancing communication or sensing capabilities [6, 7, 8]. Based on this advantage, RIS-aided ISAC systems have received extensive attention from the academic community in recent years. In [9], Wang et al. jointly optimized hybrid beamforming and RIS phase-shift matrices to improve system sensing performance. In [10], Hua et al. addressed physical layer security issues by collaboratively designing active and passive beamforming to effectively suppress information leakage while ensuring communication quality. In [11], Liu et al. discussed device-based sensing scenarios and investigated the uplink signal transmission mechanism of active targets aided by RIS.
@@ II-A Channel Model
where $\bm{u}_{i}=[0,\text{mod}(i-1,N_{1})\lambda/2,\lfloor(i-1)/N_{1}\rfloor\lambda/2]$ denotes the position of the $i$-th RIS element, $\lambda$ is the wavelength, and $i,j\in\mathcal{N}\triangleq\left\{1,2,\ldots,N\right\}$ [18].

### 16 Mar 2026 v2

@@ Abstract
Beyond-diagonal reconfigurable intelligent surfaces (BD-RIS) can realize interconnections among reflecting elements through an impedance network, thereby providing a new approach to improving the performance of integrated sensing and communication (ISAC) systems. This paper investigates the optimization problem in BD-RIS-aided multiuser ISAC system, aiming to enable a flexible trade-off design between communication and sensing performance. Specifically, we propose an optimization framework that jointly integrates multiuser interference management and sensing beam gain approximation. By jointly optimizing the precoding vector and the RIS phase-shift matrix, the proposed framework improves the multiuser communication sum rate through the interference management method and enhances the sensing performance through the beam gain approximation method. For the resulting non-convex weighted optimization problem, we employ an alternating optimization (AO) algorithm to decouple it into two subproblems, namely precoding vector optimization and phase-shift matrix optimization, where each step admits a closed-form solution. Simulation results demonstrate that the proposed BD-RIS-aided ISAC system can achieve significant improvement in the trade-offs between communication and sensing performance than the traditional diagonal RIS, verifying the effectiveness of the proposed optimization framework.
@@ I Introduction
The wireless propagation environment has a significant impact on the performance of ISAC systems. Severe occlusion in scenarios such as dense urban areas often leads to the absence of line-of-sight paths, which causes degradation in communication quality and limited sensing accuracy [5]. Reconfigurable intelligent surface (RIS), as an emerging low-cost technology, provides a feasible solution to mitigate this problem. By dynamically adjusting the phase-shifts of reflecting elements, RIS can reconfigure wireless channels, thereby enhancing communication or sensing capabilities [6, 7, 8, 9]. Based on this advantage, RIS-aided ISAC systems have received extensive attention from the academic community in recent years. In [10], Wang et al. jointly optimized hybrid beamforming and RIS phase-shift matrices to improve system sensing performance. In [11], Hua et al. addressed physical layer security issues by collaboratively designing active and passive beamforming to effectively suppress information leakage while ensuring communication quality. In [12], Liu et al. discussed device-based sensing scenarios and investigated the uplink signal transmission mechanism of active targets aided by RIS.
@@ II-A Channel Model
where $\bm{u}_{i}=[0,\text{mod}(i-1,N_{1})\lambda/2,\lfloor(i-1)/N_{1}\rfloor\lambda/2]$ denotes the position of the $i$-th RIS element, $\lambda$ is the wavelength, and $i,j\in\mathcal{N}\triangleq\left\{1,2,\ldots,N\right\}$ [19, 20].